An airplane accelerates down a runway at 3.20 m/sq.s for 32.8 seconds until is finally lifts off the ground. Determine the distance traveled before take off
Answer by Alan3354(69443)

An airplane accelerates down a runway at 3.20 m/sq.s for 32.8 seconds until is finally lifts off the ground. Determine the distance traveled before take off
------------
s = at^2/2 = 3.2*32.8^2/2
s = 1721.344 meters